use the gauss-Jordan method to solve for the following system of equations.
2x+9-z=0
3x-y+4z=1
5x+8y+3z=1

Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

2 x + 9 y + -1 z = 0
3 x + -1 y 4 z = 1
4 x + 8 y + 3 z = 1

consider equation 1 &2 Eliminate y
Multiply 1 by 1
Multiply 2 by 9
we get
2 x + 9 y + -1 z =
27 x + -9 y + 36 z =
Add the two
29 x + 0 y + 35 z =
consider equation 2 & 3 Eliminate y
Multiply 2 by 8
Multiply 3 by 1
we get
24 x + -8 y + 32 z = 8
4 x + 8 y + 3 z = 1
Add the two
28 x + 0 y + 35 z = 9
Consider (4) & (5) Eliminate x
Multiply 4 by -28
Multiply (5) by 29
we get
-812 x + -980 z = -252
812 x + 1015 z = 261
Add the two
0 x + 35 z = 9
/ 35
z = 9/35
28 x + 35 z= 9
28 x + 9 = 9
x = 0
Plug x & z in eq (1)
2 x + 9 y + -1 z = 0
0 + 9 y + - 9/35 = 0
9 y = 9/35
y = 1/35
x= 0 y= 1/35 z= 9/35

Answer by richwmiller(17219)

The other tutor didn't use the numbers supplied.
He changed 5 to 4
Is there a "y" in the first equation?
x = -34/11, y = 1, z = 31/11 if there is no y